# Improving Cramér–Rao Bound And Its Variants: An Extrinsic Geometry Perspective

### v1

@@ 4.3 Improving over Bhattacharyya/Pillai–Sinha-type bounds $\mathcal{B}^{(m)}_{\mathrm{score}}$
Fix $m\geq 1$. Let $s_{\theta}\in L^{2}(\mu)$ be the square-root embedding of the model and write $\widetilde{Y}_{j}=Y_{j}\,s_{\theta},\quad\eta_{j}=\partial_{\theta}^{j}s_{\theta},\quad j\geq 1.$ Consider the jet subspace and the score-image span
@@
By Faà di Bruno/Bell polynomials we have the jet decomposition
@@
the normal part of the remainder. Recall the normal vector
$\Pi_{m}:=P^{\perp}\eta_{m+1}\in\mathcal{T}_{m}^{\perp},$
and set the basic normal span
$\mathcal{N}_{m}=\operatorname{span}\{\Pi_{m}\}.$
@@
$\mathcal{N}_{m}^{\mathrm{aug}}\;:=\;\operatorname{span}\{\Pi_{m},\ W_{2}^{\perp},\dots,W_{m}^{\perp}\}\;\subseteq\;\mathcal{T}_{m}^{\perp}.$ (18)
(If all $W_{j}^{\perp}$ are collinear with $\Pi_{m}$ then $\mathcal{N}_{m}^{\mathrm{aug}}=\mathcal{N}_{m}$.)
@@
Lemma 6 gives $\widetilde{\mathcal{T}}_{m}\subseteq\mathcal{T}_{m}\oplus\mathcal{N}_{m}^{\mathrm{aug}}$. For any closed subspaces $V_{1}\subseteq V_{2}$ and any vector $Z$ we have $\|\operatorname{Proj}_{V_{2}}Z\|\geq\|\operatorname{Proj}_{V_{1}}Z\|$. Taking $V_{1}=\widetilde{\mathcal{T}}_{m}$ and $V_{2}=\mathcal{T}_{m}\oplus\mathcal{N}_{m}^{\mathrm{aug}}$ yields the inequality. Orthogonality of $\mathcal{T}_{m}$ and $\mathcal{N}_{m}^{\mathrm{aug}}$ gives the sum-of-squares decomposition on the left, and the equality criterion follows from standard projection geometry. ∎
Proposition 8 (Relation between basic and augmented spans; strict improvement).
With notation as above:
  1. 1.
    $\mathcal{N}_{m}^{\mathrm{aug}}=\mathcal{N}_{m}$ exactly when $W_{j}^{\perp}\in\mathcal{N}_{m}$ for every $j=2,\dots,m$. In particular, if all $W_{j}^{\perp}$ are collinear with $\Pi_{m}$, then the augmented span reduces to $\mathcal{N}_{m}$.
  2. 2.
    If $\operatorname{Proj}_{\mathcal{N}_{m}^{\mathrm{aug}}}\widetilde{Z}_{0}\neq 0$ and $\mathcal{N}_{m}^{\mathrm{aug}}\not\subseteq\widetilde{\mathcal{T}}_{m}$, then strict improvement holds:
@@
  3. 3.
    Under the condition in point 1 (so $\mathcal{N}_{m}^{\mathrm{aug}}=\mathcal{N}_{m}$), the comparison (20) reduces to
    $\mathcal{C}^{(m)}\;=\;\big\|P^{\top}\widetilde{Z}_{0}\big\|^{2}+\big\|\operatorname{Proj}_{\mathcal{N}_{m}}\widetilde{Z}_{0}\big\|^{2}\;\geq\;\mathcal{B}_{\mathrm{score}}^{(m)},$
    with strict inequality whenever $\operatorname{Proj}_{\mathcal{N}_{m}}\widetilde{Z}_{0}\neq 0$ and $\mathcal{N}_{m}\not\subseteq\widetilde{\mathcal{T}}_{m}$.
@@
(i) is immediate from the definition $\mathcal{N}_{m}^{\mathrm{aug}}=\operatorname{span}\{\Pi_{m},W_{2}^{\perp},\dots,W_{m}^{\perp}\}$.
(ii) By Proposition 7, $\mathcal{C}_{\mathrm{aug}}^{(m)}=\|\operatorname{Proj}_{\mathcal{T}_{m}\oplus\mathcal{N}_{m}^{\mathrm{aug}}}\widetilde{Z}_{0}\|^{2}\geq\|\operatorname{Proj}_{\widetilde{\mathcal{T}}_{m}}\widetilde{Z}_{0}\|^{2}$. If $\operatorname{Proj}_{\mathcal{N}_{m}^{\mathrm{aug}}}\widetilde{Z}_{0}\neq 0$ and $\mathcal{N}_{m}^{\mathrm{aug}}\not\subseteq\widetilde{\mathcal{T}}_{m}$, then the orthogonal complement $(\mathcal{T}_{m}\oplus\mathcal{N}_{m}^{\mathrm{aug}})\ominus\widetilde{\mathcal{T}}_{m}$ is nontrivial and contains a component of $\operatorname{Proj}_{\mathcal{N}_{m}^{\mathrm{aug}}}\widetilde{Z}_{0}$; therefore projection onto the larger space strictly increases the squared norm and $\mathcal{C}_{\mathrm{aug}}^{(m)}>\mathcal{B}_{\mathrm{score}}^{(m)}$.
(iii) follows by substituting $\mathcal{N}_{m}^{\mathrm{aug}}=\mathcal{N}_{m}$ into Proposition 7 and repeating the strictness argument as in (ii). ∎
@@
It is trivial to see, from the proof of Lemma 6, that we get $\mathcal{C}_{\mathrm{aug}}^{(m)}\;\geq\;\mathcal{B}_{\mathrm{score}}^{(m)}$ (and strict inequality with sufficient conditions above) with the augmented normal space being defined as $\operatorname{span}\{W_{2}^{\perp},\dots,W_{m}^{\perp}\}\;\subseteq\;\mathcal{T}_{m}^{\perp},$ without the $\Pi_{m}$ (cf. (18)). Thus we are not “cheating" by considering a higher-order derivative in the comparison.
@@ 4.4 Summary and comparison with Bhattacharyya-type bounds
The classical Bhattacharyya inequality ([8]; see also [9, 10]) generalizes the CRB by incorporating higher-order derivatives of the model density. The core idea here, as presented by Pillai and Sinha (cf. [11]), can be systematically understood within a projection framework in $L^{2}(P_{\theta})$. In their construction, one considers the span $\widetilde{\mathcal{T}}_{m}=\operatorname{span}\{\widetilde{Y}_{1},\dots,\widetilde{Y}_{m}\}$ of score derivatives (as mentioned previously, they do not exactly consider the score derivatives but the concept remains the same) and projects the centered estimator residual $\widetilde{Z}_{0}=(T-\theta)s_{\theta}$ onto $\widetilde{\mathcal{T}}_{m}$. This leads to an explicit sequence of bounds
@@
In contrast, we obtained further refinements by embedding the statistical model manifold extrinsically into the Hilbert space $L^{2}(\mu)$ and exploiting the associated differential geometry. For $m=1$, we showed that the CRB is improved by an explicit curvature correction term involving the second fundamental form $\Pi(\eta_{1},\eta_{1})$, where $\eta_{1}$ is the score direction (see Theorem 2). This yields a strictly sharper bound whenever the estimator residual has nontrivial projection in the normal direction of the model manifold – considerations of this kind are not made in the Bhattacharyya-type bounds.
For $m>1$, we constructed the jet space $\mathcal{T}_{m}=\mathrm{span}\{\eta_{1},\dots,\eta_{m}\}$ of embedded derivatives of the statistical model, and defined the augmented normal span
$\mathcal{N}^{\mathrm{aug}}_{m}=\mathrm{span}\big\{\Pi_{m},\,W_{2}^{\perp},\dots,W_{m}^{\perp}\big\},$
where $\Pi_{m}$ encodes curvature (“beyond order mm") and the $W_{j}^{\perp}$ are additional orthogonal components arising from Faà di Bruno expansions of the embedding. Our refined bound then reads
@@
$\mathcal{C}_{\mathrm{aug}}^{(m)}\;\geq\;\mathcal{B}_{\mathrm{score}}^{(m)},$
with strict inequality in generic curved models (see sufficient conditions stated in Proposition 8). Thus, the classical Bhattacharyya-type inequality/Pillai–Sinha sequence may be viewed as a score-only case, while our extrinsic geometric refinement yields systematically sharper bounds by incorporating curvature and normal components that lie beyond the score span (and one may improve upon these classical bounds even when only the curvature part is ignored; see Remark 4). These also capture the correct geometric picture of the setting under consideration. What we have shown may be succintly stated this way: Defining, for $m\geq 1$, $\mathcal{B}_{\mathrm{jet}}^{(m)}\;=\;\|\mathrm{Proj}_{\mathcal{T}_{m}}\widetilde{Z}_{0}\|^{2},$ we get the sequence of bounds $\mathcal{B}_{\mathrm{jet}}^{(1)}\leq\mathcal{B}_{\mathrm{jet}}^{(1)}\leq\mathcal{B}_{\mathrm{jet}}^{(3)}\leq\dots$ ($\mathcal{B}_{\mathrm{jet}}^{(1)}$ equalling the CRB), with
$\mathcal{B}_{\mathrm{jet}}^{(m+1)}-\mathcal{B}_{\mathrm{jet}}^{(m)}=\mathcal{C}^{(m)},$
where $\mathcal{C}^{(m)}$ is precisely the curvature correction beyond order $m$; see (16).
To conclude, our extrinsic geometry point of view results in strict improvements over classical bounds whenever the estimator error does not purely lie in the span of the score and its derivatives. This is not the perspective taken by Bhattacharyya [8]. One may argue that second-order information, as present in the curvature via second fundamental form, is taken into account in the second-order Bhattacharyya-type bounds, but this does not correspond to the curvature in any proper mathematical sense.
Some of the ideas developed in this section will be illustrated in the analytical Example 3 below.
@@ 5 Analytical examples
We illustrate Theorems 2, 5, and some of the key results stated in Section 4.3 with two fully analytical examples below; both explicitly demonstrate improvement over the CRB (i.e. $\mathcal{B}^{(1)}_{\text{score}}=\mathcal{B}^{(1)}_{\text{jet}}\equiv\mathcal{B}^{(1)}$), while the second also illuminates results comparing $\mathcal{B}_{\mathrm{score}}^{(m)}$ with our jet subspace-based curvature correction theorems.
@@
Consider the one-parameter square-root family
$s_{\theta}(x)\;=\;s_{0}(x)\,\exp\!\Big(a_{1}(\theta)\,H_{1}(x)+a_{2}(\theta)\,H_{2}(x)\Big),\qquad\theta\in\mathbb{R},$ (37)
@@
$a_{1}(\theta)=\theta,\qquad a_{2}(\theta)=\theta+\tfrac{c}{2}\,\theta^{2},$
where $c\in\mathbb{R}\setminus\{0\}$ is a fixed constant. Write $u(\theta):=\tfrac{1}{2}\log f(\cdot;\theta)=a_{1}(\theta)H_{1}+a_{2}(\theta)H_{2}+\log s_{0}$ so that
$Y_{1}=2\,\partial_{\theta}u,\qquad Y_{2}=2\,\partial_{\theta}^{2}u,\qquad\widetilde{Y}_{k}:=Y_{k}s_{\theta},\quad k=1,2.$
@@
$a_{1}^{\prime}(0)=1,\quad a_{2}^{\prime}(0)=1,\quad a_{2}^{\prime\prime}(0)=c,\qquad Y_{1}(0)=2(H_{1}+H_{2}),\quad Y_{2}(0)=2c\,H_{2}.$
@@
$\eta_{1}=\partial_{\theta}s_{\theta}=\tfrac{1}{2}Y_{1}s_{\theta},\qquad\eta_{2}=\partial_{\theta}^{2}s_{\theta}=\tfrac{1}{2}\Big(Y_{2}+\tfrac{1}{2}Y_{1}^{2}\Big)s_{\theta}.$
Evaluating at $\theta=0$ and expanding products with the Hermite product identity
$H_{m}H_{n}=\sum_{k=0}^{\min(m,n)}k!\binom{m}{k}\binom{n}{k}H_{m+n-2k}$
yields
@@
$\widetilde{\mathcal{T}}_{2}(0)=\operatorname{span}\{\widetilde{Y}_{1}(0),\widetilde{Y}_{2}(0)\}=\operatorname{span}\{2(H_{1}+H_{2})s_{0},\;2c\,H_{2}s_{0}\}=\operatorname{span}\{H_{1}s_{0},\;H_{2}s_{0}\}.$
@@
$\mathcal{T}_{2}(0)=\operatorname{span}\{\eta_{1}(0),\eta_{2}(0)\}=\operatorname{span}\Big\{(H_{1}+H_{2})s_{0},\;\big(H_{4}+2H_{3}+(5+c)H_{2}+4H_{1}+3H_{0}\big)s_{0}\Big\}.$
Thus $\mathcal{T}_{2}(0)$ is a two-dimensional subspace whose projection onto $\widetilde{\mathcal{T}}_{2}(0)$ is
$\operatorname{Proj}_{\widetilde{\mathcal{T}}_{2}(0)}\eta_{1}(0)=(H_{1}+H_{2})s_{0},\qquad\operatorname{Proj}_{\widetilde{\mathcal{T}}_{2}(0)}\eta_{2}(0)=\big(4H_{1}+(5+c)H_{2}\big)s_{0},$
which are linearly independent for all $c\neq-1$. The normal component of $\eta_{2}(0)$ relative to $\widetilde{\mathcal{T}}_{2}(0)$ is
$\operatorname{Proj}_{\widetilde{\mathcal{T}}_{2}(0)^{\perp}}\eta_{2}(0)\;=\;s_{0}\big(2H_{3}+H_{4}+3H_{0}\big),$
which is nonzero; thus $\eta_{2}(0)$ causes $\mathcal{T}_{2}(0)$ to be tilted in the directions in $\widetilde{\mathcal{T}}_{2}(0)^{\perp}$ and is precisely the geometric source of the curvature correction below.
@@
$\Gamma_{11}\;=\;\frac{\langle\eta_{2},\eta_{1}\rangle}{\langle\eta_{1},\eta_{1}\rangle}=\frac{14+2c}{3},$
@@
$\Pi_{1}\;=\;\eta_{2}-\Gamma_{11}\eta_{1}=s_{0}\Big(H_{4}+2H_{3}+\tfrac{1+c}{3}H_{2}-\tfrac{2(1+c)}{3}H_{1}+3H_{0}\Big).$
@@
$\|\Pi_{1}\|^{2}=4!\cdot 1^{2}+3!\cdot 2^{2}+2!\left(\tfrac{1+c}{3}\right)^{2}+1!\left(\tfrac{2(1+c)}{3}\right)^{2}+0!\cdot 3^{2}=57+\tfrac{2}{3}(1+c)^{2}.$
@@
which is strictly less than the full variance $\|H_{3}s_{0}\|^{2}=6$ (since $\Pi_{1}$ is not parallel to $H_{3}s_{0}$ for any real $c$). For example, at $c=1$ this evaluates to $\approx 2.41$.
At $m=2$, $\widetilde{\mathcal{T}}_{2}(0)=\operatorname{span}\{H_{1}s_{0},H_{2}s_{0}\}$ as above, so still

### v2

@@ 4.3 Improving over Bhattacharyya-type bounds $\mathcal{B}^{(m)}_{\mathrm{score}}$
Fix $m>1$. Let $s_{\theta}\in L^{2}(\mu)$ be the square root embedding of the model and write $\widetilde{Y}_{j}=Y_{j}\,s_{\theta},\quad\eta_{j}=\partial_{\theta}^{j}s_{\theta},\quad j\geq 1.$ Consider the jet subspace and the score-image span
@@
By Faà di Bruno/Bell polynomials (cf. (10), and also (11), (12), (13)) we have the jet decomposition
@@
the normal part of the remainder.
@@
$\mathcal{N}_{m}^{\mathrm{aug}}\;:=\;\operatorname{span}\{W_{2}^{\perp},\dots,W_{m}^{\perp}\}\;\subseteq\;\mathcal{T}_{m}^{\perp}.$ (18)
@@
Notice that $\mathcal{C}_{\mathrm{aug}}^{(m)}=\mathcal{B}^{(m)}_{\text{jet}}+\big\|\operatorname{Proj}_{\mathcal{N}_{m}^{\mathrm{aug}}}\widetilde{Z}_{0}\big\|^{2}$, and that one can compute $\mathcal{C}_{\mathrm{aug}}^{(m)}$ only using the $\{\eta_{k}\}_{k=1}^{m}$ offering a means to comparing bounds that can be obtained using the vectors $\{\eta_{k}\}_{k=1}^{m}$ and $\{\widetilde{Y}_{k}\}_{k=1}^{m}$. It is also obvious that $\mathcal{C}_{\mathrm{aug}}^{(m)}\geq\max\{\mathcal{B}^{(m)}_{\text{jet}},\mathcal{B}^{(m)}_{\text{score}}\}$.
@@
Lemma 6 gives $\widetilde{\mathcal{T}}_{m}\subseteq\mathcal{T}_{m}\oplus\mathcal{N}_{m}^{\mathrm{aug}}$. For any closed subspaces $V_{1}\subseteq V_{2}$ and any vector $Z$, we have $\|\operatorname{Proj}_{V_{2}}Z\|\geq\|\operatorname{Proj}_{V_{1}}Z\|$. Taking $V_{1}=\widetilde{\mathcal{T}}_{m}$ and $V_{2}=\mathcal{T}_{m}\oplus\mathcal{N}_{m}^{\mathrm{aug}}$ yields the inequality. Orthogonality of $\mathcal{T}_{m}$ and $\mathcal{N}_{m}^{\mathrm{aug}}$ gives the sum-of-squares decomposition on the left, and the equality criterion follows from standard projection geometry. ∎
Proposition 8 (Strict improvement over $\mathcal{B}_{\mathrm{score}}^{(m)}$).
With notation as above, if $\operatorname{Proj}_{\mathcal{N}_{m}^{\mathrm{aug}}}\widetilde{Z}_{0}\neq 0$ and $\mathcal{N}_{m}^{\mathrm{aug}}\not\subseteq\widetilde{\mathcal{T}}_{m}$, then strict improvement holds:
@@
By Proposition 7, $\mathcal{C}_{\mathrm{aug}}^{(m)}=\|\operatorname{Proj}_{\mathcal{T}_{m}\oplus\mathcal{N}_{m}^{\mathrm{aug}}}\widetilde{Z}_{0}\|^{2}\geq\|\operatorname{Proj}_{\widetilde{\mathcal{T}}_{m}}\widetilde{Z}_{0}\|^{2}$. If $\operatorname{Proj}_{\mathcal{N}_{m}^{\mathrm{aug}}}\widetilde{Z}_{0}\neq 0$ and $\mathcal{N}_{m}^{\mathrm{aug}}\not\subseteq\widetilde{\mathcal{T}}_{m}$, then the orthogonal complement $(\mathcal{T}_{m}\oplus\mathcal{N}_{m}^{\mathrm{aug}})\ominus\widetilde{\mathcal{T}}_{m}$ is nontrivial and contains a component of $\operatorname{Proj}_{\mathcal{N}_{m}^{\mathrm{aug}}}\widetilde{Z}_{0}$; therefore projection onto the larger space strictly increases the squared norm and $\mathcal{C}_{\mathrm{aug}}^{(m)}>\mathcal{B}_{\mathrm{score}}^{(m)}$. ∎
@@
It is worth repeating that, since the augmented normal space is defined as $\operatorname{span}\{W_{2}^{\perp},\dots,W_{m}^{\perp}\}\;\subseteq\;\mathcal{T}_{m}^{\perp},$ without the $\Pi_{m}$ (cf. (18)), we are not “cheating” by considering a higher-order derivative in the comparison by including the normal vector $\Pi_{m}$. In more formal words, Theorem 5, which is our result of the nature of Bhattacharyya, showed that considering projections onto the new (and “natural”) normal direction $\Pi_{m}$ is equivalent to incorporating a curvature-based correction, whereas Propositions 7 and 8 prove that improvement over $m$-th order Bhattacharyya-type bounds that use log-likelihood derivatives is possible using the specific structure of the $\eta_{j}$ only upto order $m$; specifically, we need to use the extrinsic directions $W_{2}^{\perp},\dots,W_{m}^{\perp}$. For the latter, it may be of interest for the reader to consider Example 4.
@@ 4.4 Summary and comparison with Bhattacharyya-type bounds
The classical Bhattacharyya inequality ([8]; see also [9, 10]) generalizes the CRB by incorporating higher-order derivatives of the model density. The core idea here can be systematically understood within a projection framework in $L^{2}(P_{\theta})$. One considers the span $\widetilde{\mathcal{T}}_{m}=\operatorname{span}\{\widetilde{Y}_{1},\dots,\widetilde{Y}_{m}\}$ of score derivatives and projects the centered estimator residual $\widetilde{Z}_{0}=(T-\theta)s_{\theta}$ onto $\widetilde{\mathcal{T}}_{m}$. This leads to an explicit sequence of bounds
@@
In contrast, we considered refinements obtained by embedding the statistical model manifold extrinsically into the Hilbert space $L^{2}(\mu)$ using the square root map and exploiting the associated differential geometry. For $m=1$, we showed that the CRB may be refined by an explicit curvature correction term involving the second fundamental form $\Pi(\eta_{1},\eta_{1})$, where $\eta_{1}$ is the first-order jet direction (see Theorem 2). This yields a strictly sharper bound whenever the estimator residual has nontrivial projection in the normal direction of the model manifold; that is, whenever we have a CRB-inefficient estimator — considerations of this kind are not made in the non-asymptotic Bhattacharyya-type bounds.
What we have shown may be succintly stated this way: Consider the jet space $\mathcal{T}_{m}=\mathrm{span}\{\eta_{1},\dots,\eta_{m}\}$ of embedded derivatives of the statistical model. Defining, for $m\geq 1$, $\mathcal{B}_{\mathrm{jet}}^{(m)}\;=\;\|\mathrm{Proj}_{\mathcal{T}_{m}}\widetilde{Z}_{0}\|^{2},$ we get the sequence of bounds $\mathcal{B}_{\mathrm{jet}}^{(1)}\leq\mathcal{B}_{\mathrm{jet}}^{(2)}\leq\mathcal{B}_{\mathrm{jet}}^{(3)}\leq\dots$ ($\mathcal{B}_{\mathrm{jet}}^{(1)}$ equalling the CRB), with
$\mathcal{B}_{\mathrm{jet}}^{(m+1)}-\mathcal{B}_{\mathrm{jet}}^{(m)}=\|\operatorname{Proj}_{\mathcal{N}_{m}}(\widetilde{Z}_{0})\|^{2},$
where $\|\operatorname{Proj}_{\mathcal{N}_{m}}(\widetilde{Z}_{0})\|^{2}$ is precisely the curvature-based correction beyond order $m$; see (16). These may be viewed as Bhattacharyya-type bounds using jets $\eta_{m}$, but the interesting part is that they come with an associated geometric picture based on curvature; also see [38] for more differential geometric intuition.
Further, we considered how one may improve the log-likelihood based bounds $\mathcal{B}_{\mathrm{score}}^{(m)}$ using only the $\eta_{j}$ upto order $m$. For $m>1$, we defined the augmented normal span
$\mathcal{N}^{\mathrm{aug}}_{m}=\mathrm{span}\big\{W_{2}^{\perp},\dots,W_{m}^{\perp}\big\},$
where the $W_{j}^{\perp}$ are orthogonal components arising from Faà di Bruno expansions of the embedding to facilitate comparing $\mathcal{T}_{m}$ with $\widetilde{\mathcal{T}}_{m}$. Our bound then reads
@@
$\mathcal{C}_{\mathrm{aug}}^{(m)}=\mathcal{B}^{(m)}_{\text{jet}}+\big\|\operatorname{Proj}_{\mathcal{N}_{m}^{\mathrm{aug}}}\widetilde{Z}_{0}\big\|^{2}\;\geq\;\mathcal{B}_{\mathrm{score}}^{(m)},$
with strict inequality in generic curved models (see sufficient conditions stated in Proposition 8). Thus, the classical Bhattacharyya-type sequence may be viewed as a score-only case, while our geometric refinement yields systematically sharper bounds by incorporating normal/extrinsic components that lie beyond the score span.
To conclude, our extrinsic geometry point of view proposes curvature-based refinements to the classical CRB whenever the estimator error does not purely lie in the span of the score. A Bhattacharyya-type extension of this was also shown, and it was argued that this intuitive geometric picture does not present itself when using the score functions associated with the (log)likelihood. Although second-order information, as present in the curvature via the second fundamental form, is taken into account in some sense in the second-order asymptotic Bhattacharyya-type bounds [17, 19], this does not translate to non-asymptotic variance bounds.
Limitations of our approach.
Having clarified our contributions, we also acknowledge the limitations. Our work largely arose from a theoretical/conceptual motivation, seeking to put forth the point that a part of the inefficiency of estimators in the sense of classical statistical bounds may be understood in terms of a refinement based on extrinsic geometry of the statistical manifold in a fixed ambient $L^{2}$ space. Some of the ideas developed here will be illustrated in the analytical examples below, which raise questions regarding ease of computation of the refinements in real situations. Our submission is that the bounds might involve more involved integral computations than when using the (log)likelihood-based scores, given the form of the $\eta_{j}$, but they can be approximated numerically in many examples of interest without much difficulty. In fact, we do just that in the asymmetric quartic location family example below rather than deriving the full set of formulae by hand, where we seek to illustrate Proposition 8; see Example 4. Our Bhattacharyya-type bound using the jets associated with $s_{\theta}$ equals
$\|\operatorname{Proj}_{\mathcal{T}_{m}}(\widetilde{Z}_{0})\|_{\theta}^{2}=b_{\eta}^{\top}(G)^{-1}b_{\eta},$
where the $i$-th component of $b_{\eta}$ is $\langle\widetilde{Z}_{0},\eta_{i}\rangle$ and $G$ is the Gram matrix associated with the $\{\eta_{i}\}_{i=1}^{m}$. Possibly, this is trickier to compute than the bounds using log-likelihood derivatives because each $\eta_{i}$ is a polynomial in the $\{Y_{j}\}_{j\leq i}$ (cf. (10)), but this complexity of the $\eta_{j}$ is also what leads to improved bounds in certain models such as the one in Example 4. Using likelihood derivatives $\frac{\partial^{i}f}{f}$ as in the original Bhattacharyya bounds is easier to handle if one has sufficient regularity such that we may differentiate under the integral with impunity. In any case, as alluded to, our main motivation is to highlight the geometric viewpoint as regards inefficiency in non-asymptotic regimes leading to our choice of fully analytical examples below that serve this purpose.
@@ 5 Analytical examples
We illustrate Theorems 2, 5, and some of the key results stated in Section 4.3 with two fully analytical examples below and one that also involves some numerics; the first two analytical examples explicitly demonstrate curvature corrections to the CRB in the event of inefficiency (i.e. $\mathcal{B}^{(1)}_{\text{score}}=\mathcal{B}^{(1)}_{\text{jet}}\equiv\mathcal{B}^{(1)}$), while the second also illuminates results comparing $\mathcal{B}_{\mathrm{score}}^{(m)}$ with our jet subspace-based curvature correction theorems. The third is specifically used to check our claim in Proposition 8 regarding the fruitfulness of incorporating alternative extrinsic directions in the normal span in addition to the $\mathcal{B}_{\mathrm{jet}}^{(m)}$ facilitating comparison and improvements over $\mathcal{B}_{\mathrm{score}}^{(m)}$.
@@
Consider the one-parameter normalized family
$f_{\theta}(x)=\frac{s_{0}(x)^{2}\exp\big(2a_{1}(\theta)H_{1}(x)+2a_{2}(\theta)H_{2}(x)\big)}{Z(\theta)},\quad s_{\theta}(x)=\sqrt{f_{\theta}(x)},$
@@
$a_{1}(\theta)=\theta,\qquad a_{2}(\theta)=\theta+\tfrac{c}{2}\theta^{2},\qquad Z(\theta)=\int_{\mathbb{R}}s_{0}(x)^{2}\exp\big(2a_{1}(\theta)H_{1}+2a_{2}(\theta)H_{2}\big)dx.$
For convergence of $Z(\theta)$ (and integrability of $f_{\theta}$), we require $2a_{2}(\theta)<1/2$. For our local computations near $\theta=0$, $c$ can then take values in a wide enough interval.
Write $u(\theta):=\tfrac{1}{2}\log f_{\theta}=a_{1}H_{1}+a_{2}H_{2}-\tfrac{1}{2}\log Z(\theta)+\log s_{0}$. Then the score functions are
$Y_{1}=2\,\partial_{\theta}u=2\big(a_{1}^{\prime}H_{1}+a_{2}^{\prime}H_{2}-\tfrac{1}{2}Z^{\prime}(\theta)/Z(\theta)\big),\quad Y_{2}=2\,\partial_{\theta}^{2}u=2\big(a_{1}^{\prime\prime}H_{1}+a_{2}^{\prime\prime}H_{2}-\tfrac{1}{2}(Z^{\prime\prime}/Z-(Z^{\prime}/Z)^{2})\big).$
@@
$a_{1}^{\prime}(0)=1,\quad a_{2}^{\prime}(0)=1,\quad a_{2}^{\prime\prime}(0)=c,\quad Z(0)=1,\quad Z^{\prime}(0)=0,\quad Z^{\prime\prime}(0)=12,$
so that the scores are
$Y_{1}(0)=2(H_{1}+H_{2}),\qquad Y_{2}(0)=2cH_{2}-12.$
@@
$\eta_{1}=\partial_{\theta}s_{\theta}=\frac{1}{2}Y_{1}s_{\theta},\qquad\eta_{2}=\partial_{\theta}^{2}s_{\theta}=\frac{1}{2}\Big(Y_{2}+\frac{1}{2}Y_{1}^{2}\Big)s_{\theta}.$
Expanding at $\theta=0$ with the Hermite product identity
$H_{m}H_{n}=\sum_{k=0}^{\min(m,n)}k!\binom{m}{k}\binom{n}{k}H_{m+n-2k},$
we obtain
@@
$\widetilde{\mathcal{T}}_{2}(0)=\operatorname{span}\{\widetilde{Y}_{1}(0),\widetilde{Y}_{2}(0)\}=\operatorname{span}\{2(H_{1}+H_{2})s_{0},\;(2c\,H_{2}-12H_{0})s_{0}\}.$
@@
$\mathcal{T}_{2}(0)=\operatorname{span}\{\eta_{1}(0),\eta_{2}(0)\}=\operatorname{span}\Big\{(H_{1}+H_{2})s_{0},\;\big(H_{4}+2H_{3}+(5+c)H_{2}+4H_{1}-3H_{0}\big)s_{0}\Big\}.$
@@
$\Gamma_{11}\;=\;\frac{\langle\eta_{2},\eta_{1}\rangle}{\langle\eta_{1},\eta_{1}\rangle},$
@@
$\Pi_{1}\;=\;\eta_{2}-\Gamma_{11}\eta_{1}=s_{0}\Big(H_{4}+2H_{3}+\tfrac{1+c}{3}H_{2}-\tfrac{2(1+c)}{3}H_{1}-3H_{0}\Big).$
@@
$\|\Pi_{1}\|^{2}=57+\tfrac{2}{3}(1+c)^{2}.$
@@
which is strictly less than the full variance $\|H_{3}s_{0}\|^{2}=6$ (since $\Pi_{1}$ is not parallel to $H_{3}s_{0}$ for any real $c$).
At $m=2$, we still get